# Towards Better Code Generation: Adaptive Decoding with Uncertainty Guidance

### v2

@@
Abstract
Code generation using large language models (LLMs) is highly sensitive to the choice of tokens during decoding, especially at points of uncertainty that critically affect the generated program’s logic. Conventional decoding methods such as greedy search and beam search apply uniform treatment to all tokens, neglecting the unique uncertainty characteristics inherent in code generation, which can result in suboptimal outputs. In this work, we conduct an empirical analysis demonstrating that a significant portion of generation errors arises from incorrect token ranking at high-uncertainty steps—where the ground truth token exists in the candidate set but fails to be ranked first.
Inspired by this insight, we introduce AdaDec, an adaptive decoding framework guided by token-level uncertainty quantified via Shannon entropy. AdaDec dynamically learns uncertainty thresholds tailored to each model and employs a pause-then-rerank mechanism with lookahead when the uncertainty surpasses these thresholds. Evaluation on the HumanEval and MBPP benchmarks reveals that AdaDec achieves up to a 15.5% improvement in Pass@1 accuracy compared to greedy decoding, matches or outperforms traditional beam search, and reduces both computational overhead and latency through targeted, selective pausing. Our findings suggest that uncertainty-aware adaptive decoding holds considerable potential for enhancing both the reliability and efficiency of code generation with LLMs.
Index Terms:
@@ I Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation [1, 2, 3, 4, 5]. Models such as Code Llama [6] and DeepSeek-Coder [7]—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts [8, 9, 10]. These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing [11, 12, 13], and test-case generation, ultimately reducing developer effort and accelerating the software development process [14, 15].
At the core of these applications lies the decoding process, which translates learned representations into executable code sequences. Among various decoding strategies, greedy decoding and beam search are widely adopted due to their simplicity and efficiency. However, these methods often suffer from early commitment to suboptimal tokens, especially under uncertainty. Such premature decisions can severely degrade the quality of generated code, given the highly structured and error-intolerant nature of programming languages.
Decoding strategies, which determine token selection, are critical to the performance of LLMs in code generation. Greedy search and beam search, standard in natural language processing, are commonly applied to code generation but remain suboptimal due to code’s unique characteristics. Unlike natural language, code exhibits distinct token uncertainty patterns. Tokens such as identifiers or line-initial keywords are highly uncertain because their selection depends on future logic. In contrast, tokens constrained by syntax or semantics, for example the keyword in within loops or repeated variable references, are more deterministic. Furthermore, code correctness is highly sensitive to decisions made at these key uncertain points. An incorrect token can drastically change program behavior; for instance, choosing “for” versus “return” at the start of a line leads to divergent control flows. Despite the critical role of token uncertainty in code generation, greedy search ignores it completely, while beam search treats all decoding steps uniformly, overlooking crucial uncertain points and reducing generation efficiency.
@@
We conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder [7] and the HumanEval dataset [16], we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding-1: The analysis indicates that many errors arise from local ranking mistakes—where the correct token is present in the candidate set but not ranked first—resulting in logic drift. To investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, computing the Shannon entropy and the rank of the ground-truth token at each decoding step. This allows us to identify and characterize potential drift points during generation. Finding-2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
@@
We introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method [17], scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete.
@@
We conduct extensive experiments on the HumanEval and MBPP benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 15.5% over Greedy decoding and achieves results that are competitive with or better than Beam Search. In terms of efficiency, AdaDec selectively pauses decoding at a low rate (between 3.95% and 11.88%), which leads to reductions in token generation of up to 54.72% and sometimes decreases overall latency. Ablation studies confirm that the learned entropy threshold effectively identifies uncertain decoding steps, and both fixed and dynamic lookahead length selection strategies improve Pass1 over Greedy decoding across different models and datasets. Finally, AdaDec demonstrates strong generalizability, consistently improving code generation accuracy across a wide range of model sizes and both benchmark datasets.
@@
  • An empirical study that reveals many code generation errors stem from local token ranking mistakes, and that Shannon entropy is a useful signal for identifying uncertain decoding steps where logic drift may occur.
  • An adaptive decoding approach named AdaDec, which uses an entropy-guided pause-then-rerank mechanism based on learned, model-specific thresholds and a lookahead strategy to improve generation quality.
  • Experimental results that show AdaDec improves Pass@1 accuracy by up to 15.5% over Greedy decoding, achieves high generation efficiency with reduced token count and latency, and generalizes effectively across diverse model sizes and benchmark datasets.
  • All code and are publicly available at [18].
II Empirical Study
@@
  • RQ1: How does token uncertainty affect the behavior of generated code?
@@
II-A Study Setup
II-A1 Models
@@
  • DeepSeek-Coder-1.3B-Instruct (DS-1.3B), DeepSeek-Coder-6.7B-Instruct (DS-6.7B) [7]: These models are trained from scratch on 2 trillion tokens comprising 87% code and 13% natural language, employing a fill-in-the-blank objective and a 16K context window to support project-level code completion and infilling tasks.
  • StableCode-Instruct-3B (ST-3B) [19]: A 2.7B parameter decoder-only model fine-tuned using Direct Preference Optimization (DPO), demonstrating state-of-the-art performance among similarly sized models on MultiPL-E and MT-Bench coding benchmarks.
  • CodeLlama-7B-Python-hf (CL-7B) [6]: A 7B parameter model specialized for Python code generation, offering capabilities such as code completion, infilling, and instruction following, and achieving strong performance on HumanEval and MBPP benchmarks.
  • Qwen3-0.6B (QW-0.6B), Qwen3-1.7B (QW-1.7B), Qwen3-4B (QW-4B), Qwen3-8B (QW-8B) [20]: A family of multilingual large language models ranging from 0.6B to 8B parameters, featuring dynamic mode switching between reasoning and fast-response modes, and supporting a wide range of tasks including code generation.
@@ II-A2 Implementation
Experiments are conducted on a machine equipped with an NVIDIA RTX 4090D GPU (24GB VRAM), an Intel(R) Xeon(R) Platinum 8474C CPU with 16 vCPUs, and 80GB of system memory.
II-B RQ1: How does token uncertainty affect the behavior of generated code?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset [16]. We then investigate the phenomenon of logic drifts during decoding, which often lead to incorrect behavior in the generated code. Finally, we explore the relationship between logic drifts and token-level uncertainty, quantified using Shannon entropy, a commonly used metric for estimating uncertainty [21].
II-B1 Method
@@
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. A mismatch is flagged when node types differ, operators or constants are inconsistent, or discrepancies occur in control structures. The mismatch is then traced back to its corresponding line in original code and extracted as a candidate drift point. This candidate is then reviewed by a human annotator to confirm whether it represents an actual drift point and to annotate the precise token where the drift begins. For example, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
@@
II-B2 Results
We identify a total of 46 drift points and 2,889 non-drift decoding steps across 46 HumanEval problems. Figure 1 presents a boxplot comparing the entropy distribution at drift points with that of all non-drift steps. The entropy values at drift points are markedly higher: the median entropy is 1.26, with a mean of 1.39, compared to a much lower median of 0.03 and mean of 0.30 for the rest. The interquartile range (IQR) for drift points is also shifted higher (0.63–2.01) relative to that of non-drift steps (0.003–0.34). The results show that tokens at drift points exhibit significantly higher entropy values compared to the rest, indicating elevated uncertainty at these critical steps.
Figure 1: Entropy comparison between drift points and non-drift decoding steps.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: This empirical pattern suggests that token uncertainty, as measured by Shannon entropy, may serve as a useful signal for identifying potential logic drifts and predicting code behavior deviations during the decoding process.
II-C RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
@@ II-C1 Method
We apply this next-token prediction procedure to a range of open-source LLMs, including four code-specific models, DS-1.3B, DS-6.7B, ST-3B, and CL-7B, as well as four general-purpose models from the Qwen3 family: QW-0.6B, QW-1.7B, QW-4B, and QW-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, 177,650 for CodeLlama, and 129,165 for Qwen3 series.
II-C2 Results
We compute the Spearman correlation between entropy and the rank of the ground-truth token across all models. As shown in Table I, the results consistently exhibit a positive correlation, further reinforcing the notion that entropy serves as a meaningful signal for identifying uncertain decoding steps. We omit p-values from the table, as the large sample sizes (approximately 130,000 to 180,000 tokens per model) yield statistically significant results with p-values effectively indistinguishable from zero.
Table I: Spearman correlation $\bm{\rho}$ between entropy and the rank of ground-truth tokens.
Model DS-1.3B DS-6.7B ST-3B CL-7B QW-0.6B QW-1.7B QW-4B QW-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.4512 0.5249 0.5028 0.4936 0.4877
Figure 2 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 3 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1, CL-7B and QW-0.6B have approximately 15% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW-1.7B) exceed a value of 4.5, one model (CL-7B) is around 2.6, and the rest fall in the range of 3.5 to 4.
Figure 2: Change in the percentage of decoding steps exceeding a given entropy threshold as the threshold increases.
Figure 3: Change in the average rank of ground-truth tokens above and below a given entropy threshold, as the threshold increases.
{mdframed}
[linecolor=gray,roundcorner=12pt,backgroundcolor=gray!15,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 2: The observed correlation between entropy and the rank of the ground-truth token suggests that entropy can be used as an indicator to adaptively pause the decoding process and rerank uncertain tokens. However, our entropy percentile analysis shows that it is difficult to define a universal, fixed entropy threshold across all models that effectively balances pause frequency and the number of reranking candidates. This highlights the need for a dynamic mechanism to determine the model-specific entropy threshold.
III Approach

### v3

@@
Abstract.
Code generation with large language models (LLMs) is highly sensitive to token selection during decoding, particularly at decision points where uncertainty strongly affects program correctness. Conventional strategies such as greedy decoding treat all tokens uniformly and fail to capture the uncertainty characteristics unique to code, often resulting in suboptimal outputs. In this work, we conduct an empirical analysis and show that a large fraction of generation errors arises from token misranking at high-uncertainty positions, where the correct token is available but not prioritized.
To address this, we introduce AdaDec, an adaptive decoding framework that employs a lookahead-based, uncertainty-aware pause-and-rerank mechanism. AdaDec automatically learns model-specific uncertainty thresholds and selectively invokes reranking when high uncertainty is detected, leveraging lookahead to refine token choice. Across HumanEval+, MBPP+, and DevEval benchmarks, AdaDec yields substantial improvements, achieving up to 20.9% absolute gains in Pass@1 accuracy compared with greedy decoding, while consistently outperforming prior adaptive decoding approaches such as AdapT. Furthermore, by applying reranking only when necessary, AdaDec reduces computational overhead and latency, enhancing efficiency alongside reliability. These findings underscore the value of uncertainty-guided decoding strategies in advancing the robustness and practicality of LLM-based code generation.
@@ 1. Introduction
Large language models (LLMs) have significantly advanced the field of code generation by learning expressive representations of programming syntax and semantics from large-scale corpora of source code and documentation (Li et al., 2023; Di and Zhang, 2025; Tian et al., 2024; Jiang et al., 2025; Lin et al., 2024). Models such as CodeLlama (Rozière et al., 2024) and DeepSeek-Coder (Guo et al., 2024)—pre-trained on billions of lines of code—exhibit strong capabilities in producing syntactically correct and context-aware code snippets from natural language prompts (Yuan et al., 2023; Sun et al., 2024; Ugare et al., 2024). These capabilities have catalyzed a range of intelligent software engineering tasks, including code completion, automated bug fixing (Fu et al., 2022; Jin et al., 2023; Xia and Zhang, 2022), and test-case generation, ultimately reducing developer effort and accelerating the software development process (Jiang et al., 2024; Huynh and Lin, 2025).
At the core of code generation with large language models (LLMs) lies the decoding process, which translates learned latent representations into discrete, executable token sequences (Chen et al., 2021a). Decoding for code generation, however, faces domain-specific challenges. Code exhibits distinctive patterns of token uncertainty: certain tokens (e.g., identifiers or control keywords at the start of a line) are intrinsically ambiguous because their correct choice depends on future logic, while others (e.g., syntactically constrained keywords or repeated identifiers) are relatively deterministic. Crucially, mistakes at a few high-uncertainty points can yield irreversible semantic errors (for example, selecting “for” versus “return” at line start fundamentally alters control flow). Figure 1 provides an example. Therefore, the choice of decoding strategy—how the next token is selected or processed—plays a decisive role in balancing efficiency, diversity, and reliability.
Figure 1. Illustration of irreversible semantic errors in code due to high-uncertainty points.
A range of strategies have been proposed: deterministic greedy decoding (Radford et al., 2019) is simple and efficient but prone to premature commitment to suboptimal tokens; sampling-based approaches such as top-$k$ (Holtzman et al., 2020) and nucleus (top-$p$) (Holtzman et al., 2020) sampling enhance diversity but often introduce syntactic or semantic errors that are unacceptable for code; and uncertainty-based methods aim to exploit token-level difficulty signals. For instance, AdapT (Zhu et al., 2023) estimates uncertainty by computing cross-entropy loss between model predictions and ground truth tokens, finding that tokens with higher loss values tend to cluster at line starts. Based on this pre-identified pattern rather than dynamic signals during decoding, AdapT assigns higher temperatures to line-initial tokens, but its reliance on randomness makes gains unstable and limited. By contrast, Uncert-CoT (Zhu et al., 2025) dynamically measures uncertainty at decoding time, employ Shannon entropy (Shannon, 1948) or probability difference (the gap between the top-1 and top-2 token probabilities) to decide whether a line-initial token is highly uncertain, in which case single-line chain-of-thought reasoning is triggered. While this avoids uniform treatment of all line starts, it overlooks uncertain tokens beyond line boundaries and has shown only marginal improvements in practice. Overall, existing methods either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, and current uncertainty-based decoding approaches have not yet explored which uncertainty signals are most appropriate to guide intervention at decoding time.
@@
To address this gap, we conduct an empirical study to better understand token-level uncertainty during the decoding process in LLM-based code generation. Using DeepSeek-Coder (Guo et al., 2024) and the HumanEval dataset (Chen et al., 2021b), we analyze decoding failures by comparing model predictions that fail test cases with their corresponding ground-truth solutions. Finding-1: The analysis reveals that many errors stem from local ranking mistakes—where the correct token exists in the candidate set but is not ranked first—thereby causing logic drift. To further investigate this phenomenon, we perform next-token prediction on 164 HumanEval problems, recording decoding information and the rank of the ground-truth token at each step. This enables us to statistically identify the locations of logic drift points and analyze the relationship between logic drift and entropy. Finding-2: The results suggest that entropy can serve as a useful signal for detecting logic drift and guiding adaptive pausing and reranking. However, due to variability across models, a fixed entropy threshold is suboptimal, highlighting the need for a dynamic, model-aware thresholding mechanism.
@@
Base on these findings, We introduce AdaDec, an uncertainty-guided adaptive decoding framework for code generation with large language models. AdaDec enhances standard decoding (e.g., greedy decoding, sampling) by incorporating a pause-then-rerank mechanism that reacts to model uncertainty, measured by Shannon entropy at each step. When uncertainty exceeds a threshold, decoding is paused and a reranking procedure is triggered. Because a fixed threshold does not generalize well across different models, AdaDec learns this threshold through a data-driven logistic regression approach. During reranking, AdaDec selects the top candidate tokens and evaluates them using a lookahead strategy inspired by A* search heuristics and a prior constrained generation method (Lu et al., 2021), scoring each candidate based on expected future continuations. The highest-scoring token is chosen for generation, and the process repeats until the sequence is complete. In contrast to prior approaches that either treat all decoding steps uniformly or rely on coarse heuristics tied to line structure, AdaDec systematically explores Shannon entropy as a token-level uncertainty signal and applies dynamic, token-level intervention across both line-initial and in-line tokens only at high-uncertainty steps, addressing the identified gap and enables targeted correction of the most consequential decision points.
@@
We conduct extensive experiments on the HumanEval+, MBPP+ and DevEval benchmarks using models ranging from 0.6B to 8B parameters. Our uncertainty-guided adaptive decoding framework, AdaDec, consistently improves Pass1 accuracy by up to 20.9% over Greedy decoding and consistently outperforms state-of-the-art adaptive decoding methods such as AdapT. In practice, AdaDec selectively pauses decoding at a low rate (averaging 7.59%), and the targeted lookahead introduces moderate additional latency, effectively balancing quality and efficiency. Ablation studies confirm that the learned entropy threshold accurately identifies uncertain decoding steps, and empirical comparisons show that using the learned threshold outperforms a fixed threshold, particularly for models where the fixed threshold deviates substantially from the learned value. Overall, AdaDec generalizes well across model scales and datasets, significantly improving code generation accuracy without incurring excessive runtime overhead.
@@
  • An empirical study that reveals many code generation errors arise from local token ranking mistakes, not only occurring at line beginnings but also sometimes within lines, and that Shannon entropy serves as a useful signal for detecting uncertain decoding steps prone to logic drift.
  • An adaptive decoding framework, AdaDec, which employs an entropy-guided pause-then-rerank mechanism with learned, model-specific thresholds and a lookahead strategy to improve generation quality.
  • Experimental results that show AdaDec improves Pass@1 accuracy by up to 20.9% over Greedy decoding, achieves high efficiency by selectively pausing decoding at a low rate, and generalizes effectively across diverse model sizes and benchmark datasets.
2. Empirical Study
@@
  • RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
@@
2.1. Study Setup
2.1.1. Models
@@
  • DeepSeek-Coder-1.3B-Instruct (DS-1.3B), DeepSeek-Coder-6.7B-Instruct (DS-6.7B) (Guo et al., 2024): These models are trained from scratch on 2 trillion tokens comprising 87% code and 13% natural language, employing a fill-in-the-blank objective and a 16K context window to support project-level code completion and infilling tasks.
  • StableCode-Instruct-3B (ST-3B) (Pinnaparaju et al., 2023): A 2.7B parameter decoder-only model fine-tuned using Direct Preference Optimization (DPO), demonstrating state-of-the-art performance among similarly sized models on MultiPL-E and MT-Bench coding benchmarks.
  • Qwen3-0.6B (QW-0.6B), Qwen3-1.7B (QW-1.7B), Qwen3-4B (QW-4B), Qwen3-8B (QW-8B) (Team, 2025): A family of multilingual large language models ranging from 0.6B to 8B parameters, featuring dynamic mode switching between reasoning and fast-response modes, and supporting a wide range of tasks including code generation.
@@ 2.1.2. Implementation
Experiments are conducted on a machine equipped with NVIDIA RTX 5880 Ada Generation GPUs (48 GB VRAM each) and an NVIDIA A100 GPU (80 GB VRAM), supported by NVIDIA Driver 550.163.01 and CUDA 12.4.
2.2. RQ1: Where do logic drifts occur in code, and how are they influenced by token uncertainty?
To address this research question, we first use DS-6.7B as a representative model to generate code for 164 programming problems in the widely used HumanEval dataset (Chen et al., 2021b). We begin by analyzing the locations of logic drifts—points during decoding where the generated sequence diverges from the correct solution—which often lead to incorrect program behavior. We then investigate how token-level uncertainty, quantified using Shannon entropy (Shannon, 1948), relates to the occurrence of such drifts.
2.2.1. Method
@@
Given the high cost of fully manual annotation, we develop a semi-automatic approach that combines abstract syntax tree (AST) comparison with human verification and fine-grained token-level annotation. To ensure annotation focuses only on faulty generations, we first filter out all generated samples that can pass the tests. Subsequently, we construct ASTs for both the ground-truth and generated code, and normalize them by removing identifiers and comments. This normalization ensures that the comparison focuses solely on structural differences. We then perform a parallel traversal to compare the two normalized ASTs node by node. A mismatch is flagged when node types differ, operators or constants are inconsistent, or discrepancies occur in control structures. The mismatch is then traced back to its corresponding line in the original code and extracted as a candidate drift point. For each mismatched line, we further determine whether the drift occurs at the beginning of the line and record this information for statistical analysis. Finally, each candidate drift point is reviewed by a human annotator to confirm whether it represents an actual drift point and to annotate the precise token where the drift begins. For example, a mismatch between the generated line if shift > len(s): and the ground-truth line if shift >= len(digits): triggers a drift point at the position of the operator >=.
@@
2.2.2. Results
We identify a total of 46 drift points and 2,889 non-drift decoding steps across 46 HumanEval problems. Statistical analysis shows that divergences at line start account for 86.96% of all drift points, while 13.04% occur within lines. This observation suggests that focusing solely on line-initial tokens for uncertainty-based intervention is insufficient, as it ignores more than 10% of logic drift points. Figure 2 presents a boxplot comparing the entropy distribution at drift points with that of all non-drift steps. For clarity, outliers are omitted in boxplots. The entropy values at drift points are markedly higher: the median entropy is 1.26, with a mean of 1.39, compared to a much lower median of 0.03 and mean of 0.30 for the rest. The interquartile range (IQR) for drift points is also shifted higher (0.63–2.01) relative to that of non-drift steps (0.003–0.34). The results show that tokens at drift points exhibit significantly higher entropy values compared to the rest, suggesting a clear relationship between entropy and logic drift.
Figure 2. Entropy comparison between drift points and non-drift decoding steps.
Finding 1: This empirical pattern indicates that focusing solely on line-start tokens for uncertainty-based intervention is insufficient, motivating the search for a suitable uncertainty signal that monitors all tokens and triggers targeted intervention at high-uncertainty steps. Subsequent analysis shows that token uncertainty measured by Shannon entropy is strongly correlated with logic drift, providing a useful signal for identifying potential drift points and predicting code behavior deviations during decoding.
2.3. RQ2: How can token uncertainty be leveraged to improve token selection during decoding?
@@ 2.3.1. Method
We apply this next-token prediction procedure to a range of open-source LLMs, including three code-specific models, DS-1.3B, DS-6.7B, and ST-3B, as well as four general-purpose models from the Qwen3 family: QW-0.6B, QW-1.7B, QW-4B, and QW-8B. For Qwen3, we disable the reasoning-specific settings (e.g., thinking mode). In total, we collect 180,879 token-level probability distributions for two DeepSeek-Coder models, 169,496 for StableCode, and 129,165 for Qwen3 series.
2.3.2. Results
We compute the Spearman correlation between entropy and the rank of the ground-truth token across all models. As shown in Table 1, the results consistently exhibit a positive correlation, further reinforcing the notion that entropy serves as a meaningful signal for identifying uncertain decoding steps. We omit p-values from the table, as the large sample sizes (approximately 130,000 to 180,000 tokens per model) yield statistically significant results with p-values effectively indistinguishable from zero.
Table 1. Spearman correlation $\bm{\rho}$ between entropy and the rank of ground-truth tokens.
Model DS-1.3B DS-6.7B ST-3B QW-0.6B QW-1.7B QW-4B QW-8B
$\bm{\rho}$ 0.4728 0.4595 0.4620 0.5249 0.5028 0.4936 0.4877
Figure 4 shows how the percentage of decoding steps with entropy exceeding a given threshold changes as the threshold increases. Figure 4 illustrates the corresponding change in the average rank of ground-truth tokens, separated by whether the entropy at each step is above or below the threshold. These plots reveal that different models exhibit varying sensitivity to entropy thresholds. For example, at a threshold of 1, QW-0.6B has approximately 15% of decoding steps exceeding the threshold, while other models remain below 10%. In terms of average rank, two models (DS-6.7B and QW-1.7B) exceed a value of 4.5, and the rest fall in the range of 3.5 to 4.
Figure 3. Change in the average rank of ground-truth tokens above and below a given entropy threshold, as the threshold increases.
Figure 4. Change in the percentage of decoding steps exceeding a given entropy threshold as the threshold increases.
Finding 2: The observed correlation between entropy and the rank of the ground-truth token suggests that entropy can be used as an indicator to adaptively pause the decoding process and rerank uncertain tokens. However, our entropy percentile analysis shows that it is difficult to define a universal, fixed entropy threshold across all models that effectively balances pause frequency and the number of reranking candidates. This highlights the need for a dynamic mechanism to determine the model-specific entropy threshold.
3. Approach